a. The probability of success on any trial of a binomial
experiment is 25%. Find the probability
that the proportion of successes in a sample of
500 is less than 22%.
b. Repeat part (a) with n = 800.
c. Repeat part (a) with n = 1,000.
1
Expert's answer
2020-12-15T08:51:21-0500

a. Find the probability of success that the proportion of success in a sample of 500 is less than 22% .

The probability of success for any trail of a binomial experiment is, p=0.25.

The sample size, n = 500

The probability of success that the proportion of success in a sample of 500 is less than 22% is,

"P(\bar{P}<0.22)=P( \frac{\bar{P}-p}{\sqrt{\frac{p(1-p)}{n}}}<\frac{0.22-0.25}{\sqrt{\frac{0.25(1-0.25)}{500}}} ) \\

=P(Z<\frac{0.22-0.25}{0.0194} \\

=p(Z<-1.55) \\

= 0.0606"

Therefore, the probability of success that the proportion of success in a sample of 500 is less than 22% is 0.0606.

b. Find the probability of success that the proportion of success in a sample of 800 is less than 22% .

The probability of success for any trail of a binomial experiment is, p=0.25.

The sample size, n=800

The probability of success that the proportion of success in a sample of 800 is less than 22% is,

"P(\bar{P}<0.22)=P( \frac{\bar{P}-p}{\sqrt{\frac{p(1-p)}{n}}}<\frac{0.22-0.25}{\sqrt{\frac{0.25(1-0.25)}{800}}} ) \\

=P(Z<\frac{0.22-0.25}{0.0153} \\

=p(Z<-1.96) \\

= 0.0250"

Therefore, the probability of success that the proportion of success in a sample of 800 is less than 22% is 0.0250.

c. Find the probability of success that the proportion of success in a sample of 1000 is less than 22%.

The probability of success for any trail of a binomial experiment is, p=0.25.

The sample size, n=1000

The probability of success that the proportion of success in a sample of 1000 is less than 22% is,

"P(\bar{P}<0.22)=P( \frac{\bar{P}-p}{\sqrt{\frac{p(1-p)}{n}}}<\frac{0.22-0.25}{\sqrt{\frac{0.25(1-0.25)}{1000}}} ) \\

=P(Z<\frac{0.22-0.25}{0.0137} \\

=p(Z<-2.19) \\

= 0.0143"

Therefore, the probability of success that the proportion of success in a sample of 1000 is less than 22% is 0.0143.